  ACPC Map

Hans Wolfgang Loidl May 14, 1993

Hagenberg Sequent Symmetry Transputer

Silicon Graphics

RISC

Linz

nCUBE 2

Salzburg

{Convex \atop C220}

Alpha Cluster Intel Hypercube iPSC/860 Sequent Balance TU-Vienna

Vienna

UNI-Vienna

UNI-Salzburg GUP Intel Hypercube iPSC/860 MasPar MP- 1101A

PPS

64 kb/s

64 kb/s 128 kb/s Transputer

MC-3DE Transpu- ter

[figure] [figure] [figure] [figure] [figure]

Vienna

Vienna

64 kb/s 64 kb/s

128 kb/s

64 kb/s

Graz